Mary has 150 nickles and dimes. She has 70 more nickles than dimes. How many of each does Mary have?
Answer by RAY100(1637)

Let n=# nickles,,,,,and d =# dimes
.
n+d=150,,,,,,,,or n =150-d
.
n=d +70,,,,,,,given
.
subst,,,,,,,(150-d) = d+70
.
150-70= 80 = 2d
.
d=40,,,,,,,,,,,,,,,n=150-40=110
.
.
check,,,,,40 +110 =150,,,,ok
.
110 = 40 +70 ,,,,,,,ok
.